OpenStudy (anonymous):

Solve for x: |3x + 2| = 2 Answer x = 0 x = 4 over 3 and x = negative 4 over 3 No Solutions x = 0 and x = negative 4 over 3

OpenStudy (katrinakaif):

3x + 2 = 2 -3x - 2 = 2 Solve for the first one : x = -0/3 = 0 Solve for the second : x = 4/-3 or simply, -4/3

hero (hero):

3x+ 2 = 2 3x+ 2 = -2 Different approach, same answer